OpenStudy (anonymous):

Give f(x)= 1/x-5 and g(x)=x^2-5 Find f(g(7)) g(f(v)) g(g(x))

OpenStudy (lgbasallote):

hint: g(7) = (7)^2 - 5 = 49 - 5 = 44 so find f(44) does that help?

OpenStudy (anonymous):

Yeah I just wanted to make sure I was right when I got 1/39

OpenStudy (lgbasallote):

seems to be right

OpenStudy (anonymous):

I'm just really confused on the other 2.

OpenStudy (lgbasallote):

just substitute v into x of f(x)

OpenStudy (lgbasallote):

for example f(v) = 1/v-5 now substitute that into the x of g(x)

OpenStudy (anonymous):

1/(v-5)^2 - 5? Should I leave it as that or go further into it?